### v2

# Patent Language Model Pretraining with ModernBERT

Abstract
Transformer-based language models such as BERT have become foundational in NLP, yet their performance degrades in specialized domains like patents, which contain long, technical, and legally structured text. Prior approaches to patent NLP have primarily relied on fine-tuning general-purpose models or domain-adapted variants pretrained with limited data. In this work, we pretrain 3 domain-specific masked language models for patents, using the ModernBERT architecture and a curated corpus of over 60 million patent records. Our approach incorporates architectural optimizations, including FlashAttention, rotary embeddings, and GLU feed-forward layers. We evaluate our models on four downstream patent classification tasks. Our model, ModernBERT-base-PT, consistently outperforms the general-purpose ModernBERT baseline on three out of four datasets and achieves competitive performance with a baseline PatentBERT. Additional experiments with ModernBERT-base-VX and Mosaic-BERT-large demonstrate that scaling the model size and customizing the tokenizer further enhance performance on selected tasks. Notably, all ModernBERT variants retain substantially faster inference—over 3× that of PatentBERT—underscoring their suitability for time-sensitive applications. These results underscore the benefits of domain-specific pretraining and architectural improvements for patent-focused NLP tasks.
Patent Language Model Pretraining with ModernBERT
Amirhossein Yousefiramandi   Ciarán Cooney Clarivate Correspondence: amirhossein.yousefiramandi@clarivate.com
1 Introduction
Since its release in 2018, BERT and other derivative encoder-based transformer models have become a mainstay of modern NLP research and applications  Devlin et al. (2019). Even within an ecosystem apparently dominated by generative Large Language Models (LLMs), BERT-based models continue to be used widely in both traditional encoder applications as well as within some LLM systems e.g., Retrieval Augmented Generation systems  Lewis et al. (2020); Wang et al. (2022).
Optimized primarily for general-domain text, BERT tends to underperform in specialized domains that feature distinct linguistic structures, such as legal or biomedical corpora Beltagy et al. (2019); Limsopatham (2021). Many such domains—including patents—exhibit idiosyncratic lexical and syntactic features, warranting domain-specific adaptation. While fine-tuning on downstream tasks remains a popular strategy, several studies have demonstrated that extending pretraining on domain-relevant corpora, or pretraining from scratch, can lead to substantial performance gains Chalkidis et al. (2020); Rasmy et al. (2021).
Patent documents are a unique blend of legal language and technical exposition, often structured and written in ways that differ sharply from general web or news text. Tasks in this domain—including classification, retrieval, and paragraph highlighting—have been addressed using both traditional ML Kamateri et al. (2022); Haghighian Roudsari et al. (2022) and transformer-based approaches Lee and Hsiang (2020); Bekamiri et al. (2024). However, the majority of BERT-based studies in this space rely on generic pretraining, which may limit their effectiveness given the domain shift.
To date, only one notable effort has reported on pretraining a BERT-style model specifically for patent documents Srebrovic and Yonamine (2020a). This indicates a significant gap in the literature, especially given the demonstrated benefits of domain adaptation in comparable fields Chalkidis et al. (2020); Qiu et al. (2021); Yan and Pei (2022). Our work aims to address this gap directly. Additionally, several technical enhancements to BERT have recently been proposed and validated, including FlashAttention  Dao et al. (2022); Shah et al. (2024), unpadding  Zeng et al. (2022), optimized MLM masking  Wettig et al. (2023), and rotary positional embeddings  Su et al. (2024). These improvements not only accelerate training, but also improve model quality and stability  Portes et al. (2023); Warner et al. (2024).
In this work, we build on these architectural and training innovations by using the ModernBERT design in conjunction with MosaicML’s Composer framework to pretrain a transformer model specifically for patent-domain tasks. Our contributions are as follows:
  • Architecture and Efficiency: We adopt the ModernBERT architecture and incorporate recent training optimizations, such as FlashAttention Dao et al. (2022), dynamic masking ratios Wettig et al. (2023), and rotary positional embeddings Su et al. (2024), as facilitated by MosaicML’s Composer framework. These enhancements significantly reduce pretraining time and compute cost compared to traditional BERT, making the model both more accessible and scalable.
  • Training Data: We construct a diverse and representative corpus that blends public patent data with proprietary patent text. The proprietary data are sections of original patent text that have been rewritten in less obscure language by subject-matter experts. This hybrid dataset enhances the model’s ability to generalize across patent categories and jurisdictions.
  • Downstream Evaluation: We benchmark our pretrained model on multiple downstream patent classification tasks and compare it against both a generic ModernBERT base model and Google’s PatentBERT Srebrovic and Yonamine (2020a). Our results demonstrate that pretraining on domain-specific patent data improves task accuracy over ModernBERT-base and leads to more efficient fine-tuning and inference in comparison with PatentBERT. Moreover, results with ModernBERT-base-VX and Mosaic-BERT-large indicate that both architectural scaling and tokenizer customization can yield further gains on selected datasets, suggesting complementary strategies for optimizing patent-domain transformers.
2 Methods
2.1 Data and Preprocessing
Pretraining a domain-specific language model with MLM loss requires a substantial volume of domain-specific data in order to compete with or exceed the performance of more generic pretrained encoder models. We construct our dataset from ~100M patent documents (mean length ~700 words), comprising original published applications and corresponding proprietary samples extracted from the Derwent World Patents Index (DWPI) [^1] https://clarivate.com/intellectual-property/patent-intelligence/derwent-world-patents-index/. This DWPI data has been constructed by subject matter experts to provide more insightful descriptions of patent content.
To prepare the data, we applied several preprocessing and quality enhancement steps. From the broader corpus, we extracted abstracts and the first independent claim from each patent, and appended DWPI titles, abstracts, and claims by matching publication numbers. Non-essential text (e.g., figure references, non-English characters) was removed. Only the first independent claim per patent was retained.
Following strategies used in FineWeb  Penedo et al. (2024b), we applied a language filter to retain only English patents and implemented text quality and repetition filters  Rae et al. (2021), using the DataTrove toolkit [^2] https://github.com/huggingface/datatrove Penedo et al. (2024b). Additional deduplication was critical given that patents are often clustered in families or legal hierarchies: we applied MinHash-based fuzzy deduplication  Penedo et al. (2024b), which computes n-gram signatures and filters out duplicates or near-duplicates. This procedure reduced the dataset from ~100M to ~64M unique patents.
Our final training dataset contains ~64M patents (~30.8B tokens), with a held-out test set of 3.4M patents (~1.6B tokens). This training corpus is approximately half the size of Google PatentBERT Srebrovic and Yonamine (2020a). Full preprocessing details are reported in Appendix  B.
2.2 Tokenization and Vocabulary Construction
To support MLM pre-training from scratch on English-language patent text, we developed a domain-specific subword tokenizer for MosaicBERT-large ( 2.2.2) using the tokenizers library Wolf et al. (2020). The tokenizer implements a Byte-Pair Encoding (BPE) model Sennrich et al. (2016) initialized with an unknown token [UNK]. All input text is normalized by Unicode NFKC normalization followed by lowercasing, reducing case-induced sparsity and consolidating visually or canonically distinct code points. Pre-tokenization uses whitespace splitting; punctuation remains attached to neighboring tokens but can be segmented by the subword model as needed.
The BPE vocabulary was trained from a streaming iterator over the pretraining corpus with a target size of 49,152 subwords and a minimum frequency threshold of 2. The following special tokens were reserved for MLM training: [UNK], [PAD], [CLS], [SEP], and [MASK]. The final vocabulary and merge rules were serialized and reused for all pretraining and evaluation. At inference time, documents are normalized, whitespace-tokenized, and mapped to subword indices; frequent technical terms and morphemes appear as single tokens, while rarer terms decompose into multiple subwords. Out-of-vocabulary character sequences are mapped to [UNK].
For comparison, we also experimented with WordPiece-style tokenization Schuster and Nakajima (2012); Devlin et al. (2019), including section-aware variants with special markers for abstracts, claims, and DWPI text Srebrovic and Yonamine (2020b). These tokenizers were trained with varying hyperparameters and evaluated using average token entropy Zouhar et al. (2023); Dagan et al. (2024) and unknown-token rates. However, preliminary end-to-end experiments showed weaker downstream performance relative to our BPE tokenizer. Since detailed results are out of scope, and all models reported in this paper use either our BPE tokenizer or ModernBERT defaults, we do not report further on these WordPiece experiments.
The other two pretrained models, ModernBERT-base-PT and ModernBERT-base-VX, employ their original publicly released tokenizers Warner et al. (2024) without modification.
Model suffixes (corpus variants).
We suffix model names to indicate the pretraining corpus variant (Table 1). -PT (clean) denotes pretraining on the Phase 2 cleaned and near-deduplicated patent corpus, which applies English language identification, repetition/quality filters, a FineWeb-style quality filter, and MinHash-based near-duplicate removal (Appendix B.2). -VX (raw) denotes pretraining on the corpus prior to Phase 2 filtering—i.e., after Phase 1 extraction/normalization and family-level deduplication only (Appendix B.1). The raw variant is larger but noisier: Phase 2 reduces tokens from $\sim 47.7$B tokens for ModernBERT-base-VX to $\sim 31.6$B tokens for ModernBERT-base-PT ($33.79\%$; Appendix B.3). See Appendix C.
Model Sfx Corpus Tokens Definition / Notes
ModernBERT-base-PT -PT Clean (Phase 2) $\sim$31.6B Phase 2 cleaned & near-deduplicated corpus: English LID; MassiveText/Gopher repetition/quality filters; FineWeb; MinHash (Appendix B.2). Phase‑2 total $\approx$31.64B (Appendix B.3).
ModernBERT-base-VX -VX Raw (pre–Phase 2) $\sim$47.7B Pre‑Phase 2 corpus (after Phase 1 & family‑level dedup only; Appendix B.1). More tokens, higher noise; aligns with $\approx$47.79B (Appendix B.3).
[…]
Table 1: Model suffix legend and corpus variants. -PT uses the Phase 2 clean corpus; -VX uses the raw corpus prior to Phase 2.
2.2.1 Model Architecture
We adopt the FlexBERT-base architecture from the ModernBERT repository [^3] https://github.com/AnswerDotAI/ModernBERT, which is a modernized BERT-based encoder tailored for efficient training Warner et al. (2024). This architecture consists of 22 encoder blocks with a hidden state size of 768 and 12 self-attention heads per layer. Gaussian Error Linear Unit (GELU) is the activation function for all feed-forward layers and Layer Normalization is applied at the beginning of each sublayer to improve training stability. The original ModernBERT has a maximum sequence length of 8192 tokens but we have limited this to 1024 for our requirements.
Following MosaicBERT Portes et al. (2023), our implementation integrates several efficiency optimizations. First, the model uses FlashAttention for the self-attention layers, enabling larger batch sizes and longer sequences Dao et al. (2022); Shah et al. (2024). Second, we adopt ALiBi (Attention with Linear Biases) positional encoding instead of learned positional embeddings Press et al. (2021). ALiBi adds a relative position bias to attention scores rather than using absolute position embeddings, which not only improves final model performance but also enables the model to generalize to sequences longer than those seen in training. Third, feed-forward layers are implemented with a Gated Linear Unit (GLU) Portes et al. (2023). This implementation sees the intermediate dense layer’s output split into two parts. One part is passed through a GELU activation and the other part is used as a gating signal (after a linear transformation) that multiplicatively modulates the activated part, enabling more flexible and expressive representations. Fourth, we scale residual connections by a factor of 0.5 following the Pre-LayerNorm setup, which stabilizes training in deeper architectures. Notably, dropout is removed entirely from the architecture without adverse effects on generalization or training stability. These model enhancements have been shown to improve training speed and stability without sacrificing model performance Portes et al. (2023).
In addition to this baseline, we pretrained two further model variants. First, a ModernBERT-base-VX configuration, which retains 22 layers and 12 attention heads but introduces adjusted intermediate dimensions and GLU expansion consistent with the Composer framework. Specifically, this variant uses a hidden size of 768, intermediate size of 1152, and GLU expansion to 2304, with RoPE $\theta$ set to 160,000 and a local attention RoPE $\theta$ of 10,000. While the total parameter count remains approximately 149M, training differed slightly from the baseline in terms of optimizer configuration and batch scheduling (see Appendix  C). This provides a controlled point of comparison against our domain-pretrained ModernBERT-base-PT.
Second, a Mosaic-BERT-large model was trained using our custom BPE tokenizer (section 2.2.2). This model scales to 28 encoder layers with a hidden size of 1024 and 16 attention heads, yielding a total of approximately 395M parameters. The feed-forward intermediate dimension is expanded to 2624, with GLU expansion to 5248. Like the base variants, it integrates FlashAttention, rotary embeddings, and Pre-LayerNorm residual scaling, but supports a maximum sequence length of 8192 tokens, making it particularly suited for patent texts with long contexts. Training followed a trapezoidal learning rate schedule with StableAdamW, distributed across 8×H100 GPUs, with dropout applied only to attention outputs (0.1). This model allows us to assess both the effect of scaling and the contribution of tokenizer customization.
Figure 1: Pretraining loss and MLM accuracy for ModernBERT-base-PT, ModernBERT-base-VX and Mosaic-BERT-large.
2.2.2 Training Procedure
Pretraining used the MLM objective. We did not include the Next Sentence Prediction task as it has been shown to provide minimal downstream performance improvement Liu et al. (2019); Izsak et al. (2021). We used a masking probability of 30% as recent findings indicate that the original probability of 15% used in BERT is suboptimal Wettig et al. (2023). Training was optimized using StableAdamW Wortsman et al. (2023), an enhancement of AdamW Loshchilov and Hutter (2018) that incorporates update clipping on a per-parameter basis similar to Adafactor Shazeer and Stern (2018). Hyperparameter details are provided in Appendix  D.
To avoid initial instability the learning rate schedule employed used a linear warmup followed by linear decay. Specifically, the learning rate was linearly ramped up from 0 to $3\times 10^{-4}$ over the first 6% of training steps during a warmup period and then linearly decayed for the remaining 94% of training steps, reaching about 2% of the peak learning rate by the end of training. We use a global batch size of 4096 sequences per step during training. This required each GPU processing a micro-batch of 128 sequences, with gradients synchronized across GPUs prior to each optimizer update. The maximum sequence length during training was set to 1024 tokens per example.
3 Results
Figure 2: Validation f1-scores for finetuning tasks across 20 epochs.
Dataset Labels Train Val. Test
WIPO 14 1731 424 533
WIPOEC 43 11349 2798 3524
HUPD 8 20772 5194 6000
DatasetCLV 5 1481 371 400
Table 2: Dataset statistics for downstream evaluation tasks.
3.1 Pretraining Evaluation
Figure 1 presents the training loss and masked accuracy for ModernBERT-base-VX, ModerBERT-base-PT and MosaicBERT-large tokenizer. The model trained with the custom BPE tokenizer exhibited a steeper decline in MLM loss in the early stages of pretraining and maintained a consistently lower loss throughout training (Figure 1). Similarly, masked token prediction accuracy rose more sharply and stabilized at a higher level compared to the WordPiece-based counterparts (Figure 1).
Pretraining results suggest that the BPE tokenizer provided a more efficient and compact representation of the patent text. BPE is known for its ability to balance vocabulary size and token granularity, particularly by effectively encoding frequent subword patterns and maintaining coherence for rare or domain-specific terms. In contrast, the standard WordPiece tokenizer—optimized for general-domain corpora—often splits complex technical terms into less informative units, making masked token prediction more difficult.
3.2 Downstream Evaluation
We validated our pretrained patent language models on downstream classification tasks using four different datasets (Table 2; Figure 3 with confidence intervals). The first two datasets come from the World International Patent Office (WIPO) [^4] https://www.wipo.int/portal/en/index.html. Another dataset is the Harvard USPTO Patent Dataset (HUPD) Suzgun et al. (2023) [^5] https://huggingface.co/datasets/HUPD/hupd. The final dataset is a proprietary dataset consisting of 5 category labels relating to data storage and networking technology. Each was split into train/validation/test sets with experiments run four times using different seeding. Reported f1 scores in Table 3 are averages of these experiments. Details on the category labels of each of the three open datasets can be found in Appendix  A.
ModernBERT-base-PT improved upon the performance of ModernBERT-base on three of the four datasets. ModernBERT-base-VX achieved broadly similar results, with slightly higher performance on the DatasetCLV dataset. Mosaic-BERT-large, trained with our custom BPE tokenizer, also demonstrated strong performance, though without consistently surpassing the smaller domain-pretrained models. PatentBERT remained strongest on HUPD, reflecting its use of USPTO data during pretraining. Taken together, these results indicate that domain-specific pretraining, architectural scaling, and tokenizer customization each provide complementary advantages for patent-specific NLP, though no single strategy yet establishes a decisive state of the art across all benchmarks.
In figure  4, evaluation f1 scores across 20 epochs of finetuing are shown (evaluation loss plots are in Appendix  E). All four plots exhibit our ModernBERT-base-PT with higher evaluation f1 scores in comparison with ModernBERT-base at all data points. This suggests there is value in domain-specific pretraininig for patent-specific downstream tasks. The plot depicting the HUPD f1 scores is noteworthy as PatentBERT outperforms the other models by a degree not seen in the other plots. We hypothize the reason for this is that USPTO data has been used during the pretraining of PatentBERT.
Model WIPO WIPOEC HUPD DatasetCLV
ModernBERT-base 0.806 0.786 0.773 0.822
ModernBERT-base-PT 0.802 0.814 0.782 0.843
ModernBERT-base-VX 0.796 0.801 0.776 0.852
MosaicBERT-large 0.787 0.796 0.772 0.850
PatentBERT 0.801 0.807 0.810 0.854
Table 3: F1 Scores for downstream classification tasks on four patent datasets.
Figure 3: Test micro-averaged F1-scores across models and datasets, reported as mean $\pm$ 95% confidence interval computed over 4 random seeds per (dataset, model). Confidence intervals are estimated using the Student’s $t$ distribution and clipped to $[0,1]$ to respect the bounded range of F1.
3.3 Finetuning Efficiency
ModernBERT has reported efficiencies in downstream training and inference Warner et al. (2024), which are further validated here. Results are reported in Table 4 and plotted with confidence intervals over 4 random seeds in Appendix E. For finetuning tasks, the average number of samples per second processed during training were 33.6, 22.7, 24.0, 9.2 and 11.5 for ModernBERT-base, ModernBERT-base-PT, ModernBERT-base-VX, MosaicBERT-large and PatentBERT, respectively. For inference, these numbers were 154.5, 157.2, 157.1, 53.3 and 44.0 samples per second. These gains stem primarily from the use of FlashAttention, ALiBi positional embeddings, and GLU-based feed-forward layers with results validating the ModernBERT architecture as a possible alternative to other BERT variants, especially in time-critical inference applications.
Model Parameters Training Inference
ModernBERT-base 149M 33.6 154.5
ModernBERT-base-PT 149M 22.7 157.2
ModernBERT-base-VX 149M 24.0 157.1
MosaicBERT-large 340M 9.2 53.3
PatentBERT 346M 11.5 44.0
Table 4: Model parameters and downstream throughput in samples per second.
4 Limitations
Language Our pretraining corpus is limited to English-language patents. Given the global nature of patent filings, expanding to multilingual corpora could significantly improve model utility in international contexts.
MLM-only objective Following Warner et al., we rely solely on the MLM objective for pretraining. While MLM has proven effective, it does not capture sentence-level semantics or inter-document relationships. Future work could explore complementary pretraining tasks such as contrastive objectives, span prediction, or retrieval-based learning tailored for patent data.
Scaling Our model was trained on a compute budget significantly smaller than that of large-scale commercial or academic models. While this demonstrates efficiency, it may limit representational capacity compared to larger-scale alternatives. Additionally, despite pretraining on  60M patents, there are potential performance increases available from further scaling of the training data.
Limited Downstream Tasks Our downstream evaluation focuses on patent classification tasks. While results are strong, they do not yet demonstrate generalization to other patent-related tasks such as similarity ranking, summarization, claim matching, or novelty detection.
5 Conclusion
In this work, we presented ModernBERT-PT, a domain-specific masked language model pretrained from scratch on a curated corpus of over 60 million patent documents. Leveraging architectural innovations such as FlashAttention, ALiBi positional embeddings, and GLU-based feed-forward layers, we demonstrated that pretraining on patent-specific data yields tangible benefits in both classification accuracy and computational efficiency. ModernBERT-base-PT outperformed a general-purpose ModernBERT baseline on three out of four downstream classification tasks and achieved competitive performance with PatentBERT, while offering substantially faster inference speeds. Additional comparisons with ModernBERT-base-VX and Mosaic-BERT-large suggest that scaling model capacity and introducing customized tokenization provide further gains on certain datasets, particularly those with longer or more technical contexts. These findings highlight the value of combining domain-specific pretraining with targeted architectural and tokenization strategies for advancing patent-domain NLP.
These findings highlight the importance of domain-specific pretraining and architecture optimization for specialized text corpora like patents. While our results validate ModernBERT-base-PT as a strong foundation for patent NLP tasks, several avenues remain for future improvement. These include scaling up the training corpus, incorporating multilingual patents, and expanding evaluations to a broader set of downstream tasks such as retrieval, summarization, and novelty detection.
With repsect to tokenization, results support growing evidence that tokenizer customization, particularly with BPE, can significantly enhance pretraining efficiency and downstream performance in domain-specific NLP. Future work could explore comparisons between BPE, SentencePiece unigram models, or byte-level tokenization schemes to further isolate the effects of subword modeling strategies in specialized corpora like patents. Overall, our work underscores the value of tailored model development for unlocking the full potential of NLP in technical and legally structured domains.
6 Acknowledgements
We would like to thank Clarivate for supporting this project and facilitating the compute resources. We would also like to thank Saurabh Mishra and Peter Keyngnaert for their support and feedback throughout the project.
Appendix A Dataset Information
WIPO Vision Dataset (WIPO)
Categories:
  • Adaptive Focus
  • Artificial Iris
  • Artificial Silicon Retina (ASR) / Retinal Prostheses
  • Augmented Reality Devices
  • Bionic Eye (System)
  • Cortical Implants
  • Drug Delivery (Vision-related)
  • Hand Wearables
  • Intraocular Lenses (IOL) with Sensors
  • Intracorneal Lenses
  • Multifocal
  • Smart Eyewear
  • Telescopic Lenses
  • Virtual Reality Devices
WIPOEC Dataset
Categories:
  • Accessories for changing body position or lifting persons – mounted in combination with a bathtub
  • Accessories for changing body position or lifting persons – mounted in combination with a toilet
  • Assistive Products for Animal Care
  • Assistive Products for Camping
  • Assistive Products for Creating Arts and Crafts
  • Assistive Products for Hunting and Fishing
  • Assistive Products for Play
  • Assistive Products for Playing and Composing Music
  • Assistive Products for Vertical Accessibility
  • Bathroom Accessories
  • Bathroom/Toilet Unit
  • Bathtub and Accessories
  • Beds and Their Accessories
  • Building Structural Components
  • Entry/Exit and Openings
  • Environment Alarms
  • Fall Detectors
  • Food Preparation
  • Golf
  • Handrails and Grab Bars
  • Laundry
  • Light Fixtures
  • Mountaineering
  • Other Furniture Accessories
  • Personal Emergency Alarm Systems and Medical Alert ID
  • Portable Travel Aids
  • Shower Unit and Accessories
  • Shower/Bathroom and Toilet Chairs
  • Sitting Arrangements and Their Accessories
  • Sports Wheelchairs
  • Storage
  • Swimming
  • Tables and Their Accessories
  • Tennis or Table Tennis
  • Toilet Seat and Accessories
  • Urinals
  • Wandering and Locating of Persons/Items
  • Wash Basin and Accessories
  • Winter Sports
  • Workplace and Domestic Safety
  • Workplace/Domestic Machinery
  • Workplace/Domestic Object Conveyance, Hoisting or Repositioning, Crane
  • Workplace/Domestic Object Securing, Gripping, Holding, Carrying and Handling
HUPD Dataset
Categories:
  • C: Chemistry; Metallurgy
  • A: Human Necessities
  • H: Electricity
  • G: Physics
  • B: Performing Operations; Transporting
  • E: Fixed Constructions
  • F: Mechanical Engineering; Lighting; Heating; Weapons; Blasting
  • D: Textiles; Paper
DatasetCLV
Categories: Five categories related to data storage and networking.
Appendix B Data Preprocessing
This appendix documents our two-phase preprocessing pipeline for large-scale patent texts used to pretrain masked language models. Phase 1 focuses on extraction, basic normalization, and family-level deduplication; Phase 2 performs FineWeb-inspired filtering and MinHash-based near-duplicate removal implemented with HuggingFace DataTrove on Databricks.
Data fields.
For each record we retain: publication number, the abstract, the first independent claim (selected from the full claim set), and the DWPI. Phase 2 is applied per field (abstract, first claim, DWPI) and results are re-assembled at the record level.
B.1 Phase 1: Extraction, Normalization, and Family-Level Deduplication
Scope. We extract fields, remove markup (HTML/XML), normalize or drop corrupted characters, strip repetitive boilerplate phrases in abstracts/claims, and remove references to figures/images. We then deduplicate by patent family: for each family we keep the most recent publication.
Scale impact (documents). From $\approx$150 M English-language patents to $\approx$70 M unique family representatives.
Algorithm 1 Phase 1: Extract, Clean, and Family-Deduplicate (pseudocode)
1:Input: Patent DB with fields $\{\texttt{pub\_no},\texttt{abstract},\texttt{claims},\texttt{DWPI},\texttt{family\_id},\texttt{pub\_date}\}$
2:Output: Clean, family-deduplicated records $\{\texttt{pub\_no},\texttt{abstract}^{*},\texttt{first\_ind\_claim}^{*},\texttt{DWPI}^{*}\}$
3:for each record $r$ do
4:  Extract pub_no, DWPI, abstract, and select first_ind_claim from claims
5:  Clean text fields: remove HTML/XML tags; normalize/drop corrupted characters
6:  Remove repetitive boilerplate prefixes; drop image/figure references
7:end for
8:Group records by family_id; within each group, keep the record with the most recent pub_date
9:return unified table with cleaned $\texttt{abstract}^{*},\texttt{first\_ind\_claim}^{*},\texttt{DWPI}^{*}$
B.2 Phase 2: FineWeb-Inspired Filtering and MinHash Near-Deduplication (Per Field)
Filtering (Phase 2a). For each field independently, we apply:
  • Language filtering via a fastText-based LID model (retain English only) Joulin et al. (2016).
  • Quality and repetition filters inspired by MassiveText/Gopher heuristics Rae et al. (2021).
  • FineWeb-style quality filter to remove subtle low-quality artifacts Penedo et al. (2024b).
[…]
Execution environment. The pipeline runs on Databricks with parallel tasks and Parquet I/O to object storage for intermediate artifacts and logs.
Algorithm 2 Phase 2: Filter and Near-Deduplicate Per Field (pseudocode)
1:Input: Phase 1 output (cleaned abstract, first_ind_claim, DWPI)
2:Output: High-quality, near-deduplicated fields
3:for each field $f\in\{\text{{abstract}},\text{{first\_ind\_claim}},\text{{DWPI}}\}$ do
4:  for each document $d$ in $f$ do
5:   if $\operatorname{LanguageID}(d)\neq\text{English}$ then
6:     discard
7:   end if
8:   if $\operatorname{RepetitionHeuristics}(d)=\text{fail}$ then
9:     discard
10:   end if
11:   if $\operatorname{QualityHeuristics}(d)=\text{fail}$ then
12:     discard
13:   end if
14:   if $\operatorname{FineWebStyleHeuristics}(d)=\text{fail}$ then
15:     discard
16:   end if
17:  end for
18:  Compute MinHash signatures on 5-gram shingles (64-bit); LSH-bucket into $14\times 8$ bands
19:  Cluster colliding documents; keep one representative per cluster; mark others as near-duplicates
20:end for
21:Re-assemble fields at the record level; count tokens before/after deduplication
B.3 Summary of Reductions
Table 5: Dataset size reductions across phases.
Phase Before After Reduction
Phase 1 (patents) $\approx$150,000,000 $\approx$70,000,000 $\approx$80,000,000
Phase 2 (tokens) 47,791,818,029 31,644,981,330 16,146,836,699   (–33.79%)
B.4 References for Filtering and Deduplication
We follow the large-scale text cleaning and deduplication practices documented in FineWeb Penedo et al. (2024b), MassiveText/Gopher Rae et al. (2021), and use fastText language identification Joulin et al. (2016). Our implementation is built on the DataTrove library Penedo et al. (2024a).
Appendix C Training Recipe
Model N tokens Epochs Train Time Training Hardware Training Strategy
ModernBERT-base-PT 31.6B 16 2.352 day 8x H100 Distributed DataParallel
ModernBERT-base-VX 47.7B 16 3.264 day 8x H100 Distributed DataParallel
MosaicBERT-large 47.7B 12 8.707 day 8x H100 Distributed DataParallel
Table 6: Training Statistics.
Appendix D Hyperparameters
Pretraining Hyperparameters
Model Optimizer LR $\beta$ $\varepsilon$ Warmup MLM
Sliding
Window (FA)
ALiBi
start size
ModernBERT-base-PT StableAdamW 3e-4 (0.90, 0.98) 1e-06 6% 30% 256 n/a
ModernBERT-base-VX StableAdamW 3e-4 (0.90, 0.98) 1e-06 6% 30% 256 n/a
MosaicBERT-large StableAdamW 2e-4 (0.90, 0.98) 1e-06 6% 30% n/a 1024
Table 7: Pretraining Hyperparameters.
Finetuning Hyperparameters
Model Optimizer LR beta weight decay epochs seeds
ModernBERT-base-PT adamw_torch_fused 3e-5 (0.9, 0.999) 3e-6 20 4
ModernBERT-base-VX adamw_torch_fused 3e-5 (0.9, 0.999) 3e-6 20 4
MosaicBERT-large adamw_torch_fused 3e-5 (0.9, 0.999) 3e-6 20 4
PatentBERT adamw_torch 2e-5 (0.9, 0.999) 1e-2 20 4
ModernBERT-base adamw_torch 5e-5 (0.9, 0.999) 5e-6 20 4
Table 8: Finetuning Hyperparameters.
Appendix E Results
Figure 4: Validation loss curves over 20 epochs for four independent fine-tuning runs.
Figure 5: Test throughput (samples per second) across models and datasets, reported as mean $\pm$ 95% confidence interval computed over 4 random seeds per (dataset, model). Confidence intervals are estimated using the Student’s $t$ distribution. Higher values indicate greater inference efficiency.

### v3

@@
 Model WIPO WIPOEC HUPD DatasetCLV
-ModernBERT-base 0.806 0.786 0.773 0.822
-ModernBERT-base-PT 0.802 0.814 0.782 0.843
-ModernBERT-base-VX 0.796 0.801 0.776 0.852
-MosaicBERT-large 0.787 0.796 0.772 0.850
-PatentBERT 0.801 0.807 0.810 0.854
-Table 3: F1 Scores for downstream classification tasks on four patent datasets.
+ModernBERT-base 0.806 $\pm$ 0.011 0.786 $\pm$ 0.005 0.773 $\pm$ 0.004 0.822 $\pm$ 0.019
+ModernBERT-base-PT 0.802 $\pm$ 0.011 0.814 $\pm$ 0.004 0.782 $\pm$ 0.003 0.843 $\pm$ 0.009
+ModernBERT-base-VX 0.796 $\pm$ 0.003 0.801 $\pm$ 0.008 0.776 $\pm$ 0.005 0.852 $\pm$ 0.005
+MosaicBERT-large 0.787 $\pm$ 0.010 0.796 $\pm$ 0.007 0.772 $\pm$ 0.006 0.850 $\pm$ 0.008
+PatentBERT 0.801 $\pm$ 0.013 0.807 $\pm$ 0.006 0.810 $\pm$ 0.001 0.854 $\pm$ 0.012
+Table 3: F1-score (mean $\pm$ std across 4 runs) for downstream classification tasks on four patent datasets.
 Figure 3: Test micro-averaged F1-scores across models and datasets, reported as mean $\pm$ 95% confidence interval computed over 4 random seeds per (dataset, model).
@@
 3.3 Finetuning Efficiency
-ModernBERT has reported efficiencies in downstream training and inference Warner et al.
-(2024), which are further validated here.
-Results are reported in Table 4 and plotted with confidence intervals over 4 random seeds in Appendix E.
-For finetuning tasks, the average number of samples per second processed during training were 33.6, 22.7, 24.0, 9.2 and 11.5 for ModernBERT-base, ModernBERT-base-PT, ModernBERT-base-VX, MosaicBERT-large and PatentBERT, respectively.
-For inference, these numbers were 154.5, 157.2, 157.1, 53.3 and 44.0 samples per second.
-These gains stem primarily from the use of FlashAttention, ALiBi positional embeddings, and GLU-based feed-forward layers with results validating the ModernBERT architecture as a possible alternative to other BERT variants, especially in time-critical inference applications.
-Model Parameters Training Inference
-ModernBERT-base 149M 33.6 154.5
-ModernBERT-base-PT 149M 22.7 157.2
-ModernBERT-base-VX 149M 24.0 157.1
-MosaicBERT-large 340M 9.2 53.3
-PatentBERT 346M 11.5 44.0
-Table 4: Model parameters and downstream throughput in samples per second.
+[…]
+Model Parameters WIPO WIPOEC HUPD DatasetCLV
+ModernBERT-base-PT 149M 83.68 $\pm$ 0.09 85.54 $\pm$ 0.15 368.42 $\pm$ 0.33 90.99 $\pm$ 0.83
+PatentBERT 346M 36.09 $\pm$ 0.06 35.50 $\pm$ 0.12 68.84 $\pm$ 0.38 35.51 $\pm$ 0.05
+ModernBERT-base 149M 84.48 $\pm$ 0.04 62.66 $\pm$ 0.52 377.41 $\pm$ 1.73 93.45 $\pm$ 0.42
+ModernBERT-base-VX 149M 84.06 $\pm$ 1.04 84.51 $\pm$ 0.38 368.09 $\pm$ 0.75 91.87 $\pm$ 0.23
+MosaicBERT-large 340M 27.44 $\pm$ 0.09 26.75 $\pm$ 0.56 127.39 $\pm$ 0.60 31.74 $\pm$ 0.30
+Table 4: Inference throughput (samples/sec; mean $\pm$ std over runs) for four patent datasets.
+Model Parameters WIPO WIPOEC HUPD DatasetCLV
+ModernBERT-base-PT 149M 21.11 $\pm$ 0.43 21.48 $\pm$ 1.01 24.98 $\pm$ 1.58 23.33 $\pm$ 0.16
+PatentBERT 346M 8.65 $\pm$ 0.14 9.29 $\pm$ 0.06 19.29 $\pm$ 0.56 8.87 $\pm$ 0.07
+ModernBERT-base 149M 21.51 $\pm$ 0.54 23.17 $\pm$ 0.93 65.04 $\pm$ 2.13 24.55 $\pm$ 0.20
+ModernBERT-base-VX 149M 20.85 $\pm$ 0.38 21.15 $\pm$ 0.80 30.38 $\pm$ 2.67 23.50 $\pm$ 0.13
+MosaicBERT-large 340M 6.19 $\pm$ 0.12 6.53 $\pm$ 0.16 17.33 $\pm$ 0.71 6.92 $\pm$ 0.07
+Table 5: Training throughput (samples/sec; mean $\pm$ std over runs) for four patent datasets.
 4 Limitations
@@
 Higher values indicate greater inference efficiency.
+Figure 6: Train throughput (samples per second) across models and datasets, reported as mean $\pm$ 95% confidence interval computed over 4 random seeds per (dataset, model).
+Confidence intervals are estimated using the Student’s $t$ distribution.
+Higher values indicate greater inference efficiency.